OpenStudy (anonymous):

A garden has width____ and length____. What is the perimeter of the garden in simplest radical form?

OpenStudy (anonymous):

\[width \sqrt{13} length 7\sqrt{13}\]

OpenStudy (anonymous):

\[a. 14\sqrt{13}units\]

OpenStudy (anonymous):

\[b. 16\sqrt{13}units \]

OpenStudy (anonymous):

c. 91 units

OpenStudy (anonymous):

\[d. 8\sqrt{13}units\]

OpenStudy (anonymous):

@HelpMe94

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

what is this?

OpenStudy (anonymous):

oh is a b c... your answers

OpenStudy (anonymous):